Solve for : log (2x + 8) = 2 Write log (2x + 8) = 2 in exponential form. Find the value of 2. > Next Question
Answer #1

Given log, (2x+8) = 2 2x+8 = x² Ans Now x²-2 .248=0 x² - 4x +2a-8=0 *(x-4)+2(x-4) = 0 (1-4)(x+2) =0 a = 4 or x = -2 Now x = -